OpenStudy (anonymous):

how would you solve: x = 8^{log_{8} 4 + log_{8} 6}

OpenStudy (anonymous):

first off \[8^{\log_8(4)}=4\]

OpenStudy (anonymous):

oh i should learn to read it is \[8^{\log_8(4)+\log_8(6)}\] right?

OpenStudy (anonymous):

yes that is correct thank you

OpenStudy (anonymous):

then you have a choice. you can write this as \[8^{\log_8(24)}=24\]

OpenStudy (anonymous):

or you can write it as \[8^{\log_8(6)}\times 8^{\log_8(4)}=6\times 4=24\] either way

OpenStudy (anonymous):

first method uses \[\log(a)+\log(b)=\log(ab)\]and the second method pretty much explains why that is true

OpenStudy (anonymous):

I was overlooking that property there, thank you!

OpenStudy (anonymous):

yw